C = 450 + 0.4Y
I = 350
G = 150
X = 70
Z = 35 + 0.1Y
T = 0.15Y
Yf = 1550
Q.2.1 Calculate the level of autonomous spending in this economy. (2)

Q.2.2 Calculate the size of the multiplier
(Note: Round your answer to two decimal places)
(4)
Q.2.3 Calculate the equilibrium level of income
(Hint: use the multiplier method)
1
Expert's answer
2020-11-30T16:34:09-0500

Q.2.1

Y = C + I + G + X – Z

Y = 450 + 0.4Y + 350 + 150 + 70 – 35 – 0.1Y

Y = 985 + 0.4(Y – T) – 0.1Y

Y = 985 + 0.4(Y – 0.15Y) – 0.1Y

Y = 985 + 0.24Y

Autonomous Spending = 985

Q.2.2

"Multiplier = \frac{1}{[1 – c(1 – t)] + m}"

In the above c is marginal propensity to consume (mpc), m is marginal propensity to import (mpm) and t is the tax rate

"Multiplier = \frac{1}{[1 – 0.4(1 – 0.15)] + 0.1} \\

Multiplier = \frac{1}{[1 – 0.4(0.85)] + 0.1} \\

Multiplier = \frac{1}{0.76} \\

Multiplier = 1.31"

Q.2.3

Y = 985 + 0.24Y

Y – 0.24Y = 985

0.76Y = 985

Equilibrium level of income = 1296